# Tri-resonant leptogenesis from modular symmetry neutrino models

### v1

@@ 6 Summary
As we know, due to the peculiarity of the observed neutrino mixing pattern, there have been intensive model-building studies on neutrino-related physics with the employment of certain discrete non-Abelian symmetries. However, a shortcoming of the conventional flavor-symmetry models is that many flavon fields need to be introduced in order to achieve a desired breaking pattern of the flavor symmetry, and the scalar potential of these flavon fields needs a complicated symmetric shaping. In recent years, the idea of modular invariance which provides substantial simplifications to these problems has been proposed and become increasingly popular.
@@
However, for some typical modular symmetry neutrino models, the right-handed neutrino mass matrix takes a form as shown in Eq. (8) which gives three degenerate right-handed neutrino masses and consequently prohibits the leptogenesis to work. But this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized. In this paper, we have considered two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one way is to modify $M_{\rm R}$ to a form as shown in Eq. (10), while another way is to consider the renormalization-group corrections for the right-handed neutrino masses.
To be specific, we have considered modular ${\rm A}_{4}$, ${\rm S}_{4}$ and ${\rm A}_{5}$ symmetry models that have a right-handed neutrino mass matrix as shown in Eq. (8). For these models, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the observed value of $\eta_{\rm B}$ can always be successfully reproduced for appropriate values of $\mu$. These results can be easily understood from the fact that given that $\mu$ is a free parameter, the resonance conditions which enhance the CP asymmetries to their maximally allowed values can always be fulfilled (this is of course helpful for us to reproduce the observed value of $\eta_{\rm B}$).
However, for the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the results will be highly restrictive. Having fixed the flavor-symmetry scale to be at the GUT scale, we reach the following conclusions for the considered representative models:
@@
  • For the modular ${\rm A}_{4}$ symmetry model given in section 3, the observed value of $\eta_{B}$ can be successfully reproduced for $M_{0}\simeq 6\times 10^{12}$ GeV ($4\times 10^{12}$ GeV or $6\times 10^{12}$ GeV) in the NO (IO) case.
  • For the modular ${\rm S}_{4}$ symmetry model given in section 4, among the five cases, only for Case E (which holds in the NO case) and $M_{0}\sim 10^{12}$ GeV can the observed value of $\eta_{\rm B}$ be successfully reproduced.
@@
  • For the modular ${\rm A}_{5}$ symmetry model given in section 5, the observed value of $\eta_{\rm B}$ can not be successfully reproduced.
@@ Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under grant Nos. 11605081, 12142507 and 12147214, the Natural Science Foundation of the Liaoning Scientific Committee under grant NO. 2022-MS-314, and the Ph. D. Research Start-up Fund of Anyang Institute of Technology (no. BSJ2021048).

### v3

@@ 7 Summary
Due to the peculiarity of the observed neutrino mixing pattern, there have been intensive model-building studies on neutrino-related physics with the employment of certain discrete non-Abelian symmetries. Particularly, in recent years, the idea of modular invariance has been proposed and become increasingly popular.
@@
However, for some typical neutrino flavor-symmetry models, the right-handed neutrino mass matrix takes a form as shown in Eq. (15) which gives three exactly degenerate right-handed neutrino masses and consequently prohibits the leptogenesis mechanism to work successfully (to be specific, there would no CP asymmetries for the decays of right-handed neutrinos). But this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized.
In this paper, we have considered four representative neutrino flavor-symmetry models which contain three exactly degenerate right-handed neutrinos, including a neutrino flavor-symmetry model realizing the TM1 mixing, the modular ${\rm A}_{4}$ symmetry model, the modular ${\rm S}_{4}$ symmetry model, and the modular ${\rm A}_{5}$ symmetry model. For these models, we have performed a careful numerical calculation of the RGE induced mass splittings of right-handed neutrinos by employing the formulas as described in section 2.2 and subsequently the resulting baryon asymmetry by employing the formulas as described in section 2.1. The results show that, for appropriate values of the free parameters $M_{0}$ and $\tan\beta$, the observed value of $\eta_{\rm B}$ has chance to be successfully reproduced. To be specific, we reach the following conclusions for the considered representative models:
  • For the TM1 model in section 3, the observed value of $\eta_{\rm B}$ can be successfully reproduced for $M_{0}\sim 7\times 10^{12}$ GeV.
@@
  • For the modular ${\rm A}_{4}$ symmetry model in section 4, in the NO case the observed value of $\eta_{\rm B}$ can be successfully reproduced when $M_{0}$ is approximately between $10^{12}$ GeV and $10^{13}$ GeV. While in the IO case the observed value of $\eta_{\rm B}$ can not be reproduced successfully since the magnitudes of CP asymmetries are too large.
  • For the modular ${\rm S}_{4}$ symmetry model in section 5, for Case A and Case C there is an upper bound on $\tan\beta$ (about 45 and 28, respectively) and for each allowed value of $\tan\beta$ there may exist two allowed values of $M_{0}$. Importantly, Case C allows relatively low values of $M_{0}$, which may be reachable for future direct measurements. For the other cases, the allowed values of $M_{0}$ (about $10^{14}$, $10^{9}$ and $10^{13}$ GeV for Case B, Case D and Case E, respectively) are almost independent of $\tan\beta$.
@@
  • For the modular ${\rm A}_{5}$ symmetry model in section 6, the observed value of $\eta_{\rm B}$ can be successfully reproduced for $M_{0}$ in the range between $10^{7}$ and $10^{10}$ GeV for $\tan\beta$ from 10 to 50.
For the four models discussed above, the energy scales required for successful leptogenesis are relatively high. This appears to contradict the common expectation that the resonant leptogenesis mechanism can lower the energy scale for successful leptogenesis to around the TeV scale. The reason lies in the distinct parameter constraints between general models and the ones studied here: In typical models, parameters are either unconstrained or subject to only loose restrictions, allowing resonant leptogenesis to effectively reduce the required energy scale to TeV or nearby values for successful leptogenesis. In contrast, the flavor structures of the models considered in this paper are tightly constrained by the imposed flavor symmetries which result in fewer free parameters (in fact, we are only left with the right-handed neutrino mass scale and $\tan\beta$ as free parameters) and thus stronger predictive power. This strong constraint has critical implications for the specific scenario of leptogenesis triggered by renormalization group running explored in our work: For these models, the TeV scale no longer accommodates successful leptogenesis. In fact, for some of them (see, e.g., Figures 9 and 12), a TeV-scale setup would lead to a lepton asymmetry larger than the observed value.
@@ Acknowledgments
Zhi-Cheng Liu acknowledges the support from the Ph. D. Research Start-up Fund of Anyang Institute of Technology (No. BSJ2021048). Zhen-hua Zhao acknowledges the support from the National Natural Science Foundation of China under Grant No. 12475112, Liaoning Revitalization Talents Program under Grant No. XLYC2403152, and the Basic Research Business Fees for Universities in Liaoning Province under Grant No. LJ212410165050.